Grade 12Physical Chemistry

a vessel has Nitrogen gas and water vapours in equilibrium with liquid water at a total pressure of 1 ATM. the partial pressure of water vapours is 0.3 ATM. the volume of this vessel is reduced to one third of the original volume at same temperature then total pressure of the system is

Akash Sharma
10 Years agoGrade 12

1 Answer

Shikhar Panchal
5 Years ago
We are given that P(net) = 1 atm
Hence , p(H20) + p(N2) = 1 atm
P(H20) = 0.3 atm
Let volume of the old vessel be V
And volume of new vessel be V/3
But water is incompressible 
So we know that p(H20) = constant = 0.3 atm
Therefore , by the following relation we can get our answer, 
P is inversely proportional to volume (Boyles law)
We can say that if v decreases by 3 times then p(N2) increases by 3 times
Hence, P(net) = 3*P(N2) + p(H2O) = 2.1 + 0.3 = 2.4 atm